The height of a poplar tree in feet at age t years can be modeled by the function h(t) = 6+3ln(t+1). Use the model to predict the number of years when the height will exceed 18 feet.
Answer by Alan3354(69443)

The height of a poplar tree in feet at age t years can be modeled by the function h(t) = 6+3ln(t+1). Use the model to predict the number of years when the height will exceed 18 feet.
---------------------
Find when it is 18 feet:
h(t) = 6+3ln(t+1) = 18
3ln(t+1) = 12
ln(t+1) = 4
t+1 = e^4
t > e^4 - 1 years